# Augmented doubly robust post-imputation inference for proteomic data

### v1

@@ 1 Introduction
Recently single-cell RNA sequencing technology has fueled a revolution in our ability to study biological processes. However, mRNA transcript abundances are only a weakly correlated precursor to protein abundances (Vogel and Marcotte, 2012; Liu et al., 2016; Tasaki et al., 2022). And it is the protein that carries out the more fundamental roles of molecular mechanisms in cellular processes. Developments in mass spectrometry proteomic technology have greatly enhanced the quantitative analysis of proteins related to human health and disease. Nevertheless, such analyses often encounter challenges due to a high rate of missingness, especially for single-cell data, resulting from various technological factors (Vanderaa and Gatto, 2023). While missingness significantly impacts the validity and efficiency of downstream tasks, the optimal method for handling missing data in proteomics remains a subject of active debate, and is an area in need of novel methodological advancements (Shen et al., 2022).
@@
Currently, a major focus of discussion in the field is on the choice of imputation method (Vanderaa and Gatto, 2023; Wei et al., 2018), which is used to infer peptide abundance, a subunit of proteins. It is common practice that imputed values are directly plugged into the original dataset, followed by downstream analyses as if the imputed values were the original observed ones (“Plugin method”). With this method, the assumption is that the imputed data accurately represents the original data. Therefore, the precision of the imputed result is crucial for a valid downstream analysis. A substantial ongoing research effort is to search and experiment with numerous imputation methods to determine the optimal one, including sample matching methods (Stuart and Satija, 2019), matrix factorization methods (Hastie et al., 2015), deep learning methods (Yoon et al., 2018; Qiu et al., 2020; Du et al., 2022) and more (Wang et al., 2016; Chen et al., 2017). See Harris et al. (2023); Välikangas et al. (2018); Liu and Dongre (2021) for a comprehensive review.
@@
Most of the aforementioned methods rely on a high-dimensionality and robust intercorrelation structure of the measured peptides. Such characteristics of proteomic data provide a solid foundation for various imputation algorithms; however, this approach may not be ideal when the downstream analysis plan is based on the Plugin method. There are two reasons for this. First, the aim of retrieving the original outcomes via imputation is not optimal in some downstream analyses. Consider a linear model in which we regress each peptide abundance in some low-dimensional covariates. In this context, the optimal choice for imputation is the conditional mean abundance based on these covariates. When the Plugin method is combined with high-dimensional imputation models, we are attempting to get closer to the original outcome, rather than the conditional mean, which may introduce additional variance into estimated regression coefficients. Second, when the full high-dimensional dataset is used for imputation, a systematic bias can be introduced into the imputed data, causing false discovery due to confounding. A recent paper by Agarwal et al. (2020) investigates this issue using transcriptomic datasets. They show that if the dataset contains a number of differentially expressed genes, a naive application of the Plugin method results in notably inflated False Discovery Rates (FDR). This inflation does not occur when none of the genes are differentially expressed, which indicates that the source of the FDR inflation is the cross-use of high-dimensional data for imputation. More discussions on this can be found in Andrews and Hemberg (2018); Ly and Vingron (2022). Similar post-imputation inference issues remain for proteomic studies.
@@
One approach that can circumvent these issues is to use only complete data for analysis and simply ignore missingness (“Complete method”). This provides a simple and valid way to prevent problems from imputation under certain missingness assumptions. However, it discards any indirect information on missing outcomes and is especially vulnerable to low power with small sample sizes. Multiple imputation (Rubin, 1987) is another possible approach, which provides a general framework for obtaining valid statistical inferences while incorporating the imputation uncertainty. This technique avoids denoising and involves generating multiple complete datasets by filling in missing data with several plausible imputations. The resulting test statistic incorporates variances both within and between datasets to compute the total variance. There have been some attempts to apply this framework to proteomic data (Yin et al., 2016; Gianetto et al., 2020). Some noticeable challenges involved in using this approach include its empirical conservativeness (Chion et al., 2022), computational burden (Brini and van den Heuvel, 2023), and the lack of a straightforward expression for test statistics (Meng, 1994).
@@
In this paper, we propose an alternative framework motivated by doubly robust estimation (Scharfstein et al., 1999), a widely used procedure to estimate mean outcomes. Our purpose is to establish a valid and efficient inference framework that is well-harmonized with high-dimensional imputation models. Estimating mean outcomes is a significant area of research, especially when certain outcomes are not observable and a propensity score (probability of observation) depends on measured covariates. Then observed outcomes do not accurately represent the entire population due to the covariate mismatch. Therefore, instead of simply averaging the observed outcomes, one first constructs an outcome model by regressing the outcomes on covariates related to the propensity score and averaging the fitted values over the entire population. A doubly robust estimator incorporates an additional term to correct for the first-order bias of the fitted outcomes. While two nuisance estimators – an outcome estimator and a propensity score estimator – are employed, this approach enjoys a “double robustness” property, which means that the statistic remains consistent as long as at least one of the nuisance estimators is consistent (Robins and Rotnitzky, 1995). Several recent papers extend this strategy to estimation problems beyond the mean outcome (Kennedy, 2023; Fisher and Fisher, 2023; Díaz et al., 2018; Qiu and Messer, 2023). In particular, Kennedy (2023) uses each summand of the doubly robust estimator as a pseudo-outcome to measure a conditional average treatment effect in a nonparametric regression setting.
@@
In our framework, the propensity score is estimated through a conventional logit model to enjoy a fast rate of parametric convergence, while the outcome model is estimated through a flexible machine-learning method that can handle high-dimensional variables and their complex relationships. Our framework not only calls for, but also deliberately invites powerful modern methods because it includes a built-in mechanism to push the estimator towards achieving $\sqrt{n}$-consistency, even when the employed imputation method fails to achieve a sufficiently fast rate. In our simulations and data study, we use a variant of VAE models called VAEIT (Du et al., 2022) to fit the outcomes; see Appendix B for more details. The VAEIT model utilizes both low-dimensional covariates and high-dimensional proteomic data, and offers enough flexibility to handle missing data as well as non-linear dependency.
@@
In high-dimensional nuisance parameter estimation, Jiang et al. (2022) and Yadlowsky (2022) derived consistency results for estimated conditional treatment effect with sparsity or distributional assumptions. Double machine learning, proposed by Chernozhukov et al. (2018), provides a framework for building an efficient estimator of low-dimensional parameters, with nuisance functions estimated using a high-dimensional black-box model. More papers based on semiparametric nuisance estimation are summarized in Davidian (2022). Most of the aforementioned references use the same set of high-dimensional variables for both nuisance functions. Other lines of investigation, including Berrevoets et al. (2023), Zhao and Ding (2022), and Little et al. (2012), explore an estimator for average treatment effect when some data are missing. They measure the effect size by adjusting the covariate distributions of treatment and control groups separately and computing the outcome difference.
@@ 2.1 Background
If the rate of missingness is only related to a measured covariate ($C_{i}\perp Y_{i}|\mathbf{W}_{i}$), a simple strategy of excluding missing samples provides valid inferential results (Little, 1992); but it comes at the expense of a reduced sample size. Therefore, we consider the pseudo-outcome approach, which can offer better statistical efficiency. In an ideal scenario, when the conditional mean $\mathbb{E}[Y_{i}\mid\mathbf{W}_{i}]$ is available, replacing the outcome data with this value will provide valid and efficient inference. This approach has been explored in causal inference studies. While the typical average treatment effect estimates $\mathbb{E}[Y_{i}]$, the conditional average treatment effect seeks an individualized conditional outcome $\mathbb{E}[Y_{i}\mid\mathbf{W}_{i}]$, especially when $Y_{i}$ is not observable for counterfactual cases. Several recent papers address this issue by utilizing pseudo-outcomes, which have the same conditional means as the original outcomes, and fitting a regression against them as if they were observed data (Kennedy, 2023; Fisher and Fisher, 2023; Semenova and Chernozhukov, 2021; Díaz et al., 2018). This approach allows for achieving desirable properties such as robustness and efficiency through a selection of appropriate pseudo-outcomes.
Inspired by these studies, we further extend the pseudo-outcome framework in linear regression. We especially focus on the doubly robust estimator suggested by Scharfstein et al. (1999), which is extensively used to estimate the mean outcome $\mathbb{E}[Y_{i}]$. This estimator is defined upon the construction of two nuisance functions;
@@
with estimated outcome/propensity model $\hat{\mu}$ and $\hat{\delta}$. Appealing properties of the estimator arise from its second-order nuisance estimation error, that is,
@@ 2.2 An augmented doubly robust estimator $\hat{\bm{\beta}}_{UW}$
Under the same conditions in Theorem 2.2, further assume that $\|(1-{\delta_{i}}/{\hat{\delta_{i}}})(\hat{\nu_{i}}-\nu_{i})\|_{2}=o_{\mathbb{P}}(n^{-1/2})$. Then the estimator $\hat{\bm{\beta}}_{UW}$ defined in (4) is asymptotically normal:
@@
$\hat{\Sigma}_{UW}=\left(\frac{1}{n}\sum_{i=1}^{n}\mathbf{W}_{i}\mathbf{W}_{i}^{T}\right)^{-1}\left(\frac{1}{n}\sum_{i=1}^{n}(\hat{Y}_{i}^{UW}-\mathbf{W}_{i}\hat{\bm{\beta}}_{UW})^{2}\mathbf{W}_{i}\mathbf{W}_{i}^{T}\right)\left(\frac{1}{n}\sum_{i=1}^{n}\mathbf{W}_{i}\mathbf{W}_{i}^{T}\right)^{-1}$ (5)
@@ 3.1 The input data and the algorithm
The observed abundance of $p$ peptides from $n$ samples can be written as an $n\times p$ matrix $\mathbf{C}\odot\mathbf{Y}\in\mathbb{R}^{n\times p}$, where $\mathbf{C}\in\{0,1\}^{n\times p}$ indicates the entry-wise missingness and $\mathbf{Y}\in\mathbb{R}^{n\times p}$ is the full data matrix without missing. Here $\odot$” stands for the component-wise product. Only $\mathbf{C}$ and $\mathbf{C}\odot\mathbf{Y}$ are available. Also observed is a covariate data matrix $\mathbf{W}\in\mathbb{R}^{n\times q}$. The inference task is to test the significance of regression coefficients for the low dimensional covariates in $\mathbf{W}$ on each peptide. To this end, we will obtain individual p-values for each peptide using the asymptotic results presented in the previous section, and then apply a multiple testing framework such as the Benjamini-Hochberg procedure (Benjamini and Hochberg, 1995).
This procedure involves regressing each column of $\mathbf{Y}$ on both the low-dimensional covariate $\mathbf{W}$ and the other peptides as the high-dimensional auxiliary covariate $\mathbf{U}$. The fitted regression function is used as $\hat{\nu}$ in the imputation method described in Section 2. An additional challenge is that each column of $\mathbf{Y}$ has many missing entries, even when used as a covariate in the regression problem. To address this issue, we use variational autoencoder, a deep neural network tool that allows for flexible input and simultaneous estimation of the multi-response regression. The algorithm used in our simulation and data examples is VAEIT (Du et al., 2022, see Appendix B for details). The whole procedure is summarized in Algorithm 1.

### v2

@@ 1 Introduction
Recently single-cell RNA sequencing technology has fueled a revolution in our ability to study biological processes. However, mRNA transcript abundances are only a weakly correlated precursor to protein abundances (Vogel and Marcotte, 2012; Liu, Beyer and Aebersold, 2016; Tasaki et al., 2022). And it is the protein that carries out the more fundamental roles of molecular mechanisms in cellular processes. Developments in mass spectrometry proteomic technology have greatly enhanced the quantitative analysis of proteins related to human health and disease. Nevertheless, such analyses often encounter challenges due to a high rate of missingness, especially for single-cell data, resulting from various technological factors (Vanderaa and Gatto, 2023). While missingness significantly impacts the validity and efficiency of downstream tasks, the optimal method for handling missing data in proteomics remains a subject of active debate, and is an area in need of novel methodological advancements (Shen et al., 2022).
To assess protein abundance, the measured units are peptides, which are subunits of a protein. Peptides that are present in the sample matrix, but not assigned an abundance value for the observations in a batch are considered missing. Missingness can be attributed to a variety of technical factors that lead to a failure to measure abundance across all observed spectra (Webb-Robertson et al., 2015; Bramer et al., 2021; Vanderaa and Gatto, 2023). Missing patterns have been reported to be close to random (Brenes et al., 2019), meaning that while the propensity depends weakly on measured covariates, it tends not to depend on the true value of the abundance. Peptides with very low abundance are more likely to be missing; however, by design, mass spectrometry is calibrated to measure common peptides and hence no measurements are made on the majority of low abundance peptides. Thus while some missing values are missing not at random (MNAR), the vast majority follow the missing at random (MAR) assumption.
@@
Currently, a major focus of discussion in the field is on the choice of imputation method (Vanderaa and Gatto, 2023; Wei et al., 2018), which is used to infer peptide abundance. It is common practice that imputed values are directly plugged into the original dataset, followed by downstream analyses as if the imputed values were the original observed ones (“Plugin method”). With this method, the assumption is that the imputed data accurately represents the original data. Therefore, the precision of the imputed result is crucial for a valid downstream analysis. A substantial ongoing research effort is to search and experiment with numerous imputation methods to determine the optimal one, including sample matching methods (Stuart and Satija, 2019), matrix factorization methods (Hastie et al., 2015), deep learning methods (Yoon, Jordon and Schaar, 2018; Qiu, Zheng and Gevaert, 2020; Du et al., 2022) and more (Wang et al., 2016; Chen et al., 2017). See Harris et al. (2023); Välikangas, Suomi and Elo (2018); Liu and Dongre (2021) for a comprehensive review.
@@
Most of the aforementioned methods rely on a high-dimensionality and robust intercorrelation structure of the measured peptides. Such characteristics of proteomic data provide a solid foundation for various imputation algorithms; however, this approach may not be ideal when the downstream analysis plan is based on the Plugin method. There are two reasons for this. First, the aim of retrieving the original outcomes via imputation is not optimal in some downstream analyses. Consider a linear model in which we regress each peptide abundance in some low-dimensional covariates. In this context, the optimal choice for imputation is the conditional mean abundance based on these covariates. When the Plugin method is combined with high-dimensional imputation models, we are attempting to get closer to the original outcome, rather than the conditional mean, which may introduce additional variance into estimated regression coefficients. Second, when the full high-dimensional dataset is used for imputation, a systematic bias can be introduced into the imputed data, causing false discovery due to confounding. A recent paper by Agarwal, Wang and Zhang (2020) investigates this issue using transcriptomic datasets. They show that if the dataset contains a number of differentially expressed genes, a naive application of the Plugin method results in notably inflated False Discovery Rates (FDR). This inflation does not occur when none of the genes are differentially expressed, which indicates that the source of the FDR inflation is the cross-use of high-dimensional data for imputation. More discussions on this can be found in Andrews and Hemberg (2018); Ly and Vingron (2022). Similar post-imputation inference issues remain for proteomic studies.
@@
One approach that can circumvent these issues is to use only complete data for analysis and simply ignore missingness (“Complete method”). This provides a simple and valid way to prevent problems from imputation under certain missingness assumptions. However, it discards any indirect information on missing outcomes and is especially vulnerable to low power with small sample sizes. Multiple imputation (Rubin, 1987) is another possible approach, which provides a general framework for obtaining valid statistical inferences while incorporating the imputation uncertainty. This technique avoids denoising and involves generating multiple complete datasets by filling in missing data with several plausible imputations. The resulting test statistic incorporates variances both within and between datasets to compute the total variance. There have been some attempts to apply this framework to proteomic data (Yin et al., 2016; Gianetto et al., 2020). Some noticeable challenges involved in using this approach include its empirical conservativeness (Chion, Carapito and Bertrand, 2022), computational burden (Brini and van den Heuvel, 2023), and the lack of a straightforward expression for test statistics (Meng, 1994).
@@
In this paper, we propose an alternative framework motivated by doubly robust estimation (Scharfstein, Rotnitzky and Robins, 1999), a widely used procedure to estimate mean outcomes. Our purpose is to establish a valid and efficient inference framework that is well-harmonized with high-dimensional imputation models. Estimating mean outcomes is a significant area of research, especially when certain outcomes are not observable and a propensity score (probability of observation) depends on measured covariates. Then observed outcomes do not accurately represent the entire population due to the covariate mismatch. Therefore, instead of simply averaging the observed outcomes, one first constructs an outcome model by regressing the outcomes on covariates related to the propensity score and averaging the fitted values over the entire population. A doubly robust estimator incorporates an additional term to correct for the first-order bias of the fitted outcomes. While two nuisance estimators – an outcome estimator and a propensity score estimator – are employed, this approach enjoys a “double robustness” property, which means that the statistic remains consistent as long as at least one of the nuisance estimators is consistent (Robins and Rotnitzky, 1995). Several recent papers extend this strategy to estimation problems beyond the mean outcome (Kennedy, 2023; Fisher and Fisher, 2023; Díaz, Savenkov and Ballman, 2018; Qiu and Messer, 2023). In particular, Kennedy (2023) uses each summand of the doubly robust estimator as a pseudo-outcome to measure a conditional average treatment effect in a nonparametric regression setting.
@@
In our framework, the propensity score is estimated through a conventional logit model to enjoy a fast rate of parametric convergence, while the outcome model is estimated through a flexible machine-learning method that can handle high-dimensional variables and their complex relationships. Our framework not only calls for, but also deliberately invites powerful modern methods because it includes a built-in mechanism to push the estimator towards achieving $\sqrt{n}$-consistency, even when the employed imputation method fails to achieve a sufficiently fast rate. In our simulations and data study, we use a variant of VAE models called VAEIT (Du et al., 2022) to fit the outcomes; see B for more details. The VAEIT model utilizes both low-dimensional covariates and high-dimensional proteomic data, and offers enough flexibility to handle missing data as well as non-linear dependency.
@@
In high-dimensional nuisance parameter estimation, Jiang et al. (2022) and Yadlowsky (2022) derived consistency results for estimated conditional treatment effect with sparsity or distributional assumptions. Double machine learning, proposed by Chernozhukov et al. (2018), provides a framework for building an efficient estimator of low-dimensional parameters, with nuisance functions estimated using a high-dimensional black-box model. More papers based on semiparametric nuisance estimation are summarized in Davidian (2022). Most of the aforementioned references use the same set of high-dimensional variables for both nuisance functions. Other lines of investigation, including Berrevoets et al. (2023), Zhao and Ding (2022), and Little et al. (2012), explore an estimator for average treatment effect when some data are missing. They measure the effect size by adjusting the covariate distributions of treatment and control groups separately and computing the outcome difference. In a matrix completion problem, Chen et al. (2019) suggests a method for denoising a matrix with corrupted and missing entries based on low-rank decomposition. Their method debiases the initial rank-constrained estimator and provides confidence intervals with distributional guarantees. Shao and Zhang (2023), Gui, Barber and Ma (2023) suggests entrywise predictive inferences using a conformal prediction framework. This line of work aims to provide uncertainty quantification for each imputed entry and doesn’t consider the associations between the outcomes and the treatment, which is substantially different from our work.
@@ 2.1 Background
If the rate of missingness is only related to a measured covariate ($C_{i}\perp Y_{i}|\mathbf{W}_{i}$), a simple strategy of excluding missing samples provides valid inferential results (Little, 1992); but it comes at the expense of a reduced sample size. Therefore, we consider the pseudo-outcome approach, which can offer better statistical efficiency. In an ideal scenario, when the conditional mean $\mathbb{E}[Y_{i}\mid\mathbf{W}_{i}]$ is available, replacing the outcome data with this value will provide valid and efficient inference. This approach has been explored in causal inference studies. While the typical average treatment effect estimates $\mathbb{E}[Y_{i}]$, the conditional average treatment effect seeks an individualized conditional outcome $\mathbb{E}[Y_{i}\mid\mathbf{W}_{i}]$, especially when $Y_{i}$ is not observable for counterfactual cases. Several recent papers address this issue by utilizing pseudo-outcomes, which have the same conditional means as the original outcomes, and fitting a regression against them as if they were observed data (Kennedy, 2023; Fisher and Fisher, 2023; Semenova and Chernozhukov, 2021; Díaz, Savenkov and Ballman, 2018). This approach allows for achieving desirable properties such as robustness and efficiency through a selection of appropriate pseudo-outcomes.
Inspired by these studies, we further extend the pseudo-outcome framework in linear regression. We especially focus on the doubly robust estimator suggested by Scharfstein, Rotnitzky and Robins (1999), which is extensively used to estimate the mean outcome $\mathbb{E}[Y_{i}]$. This estimator is defined upon the construction of two nuisance functions;
@@
with estimated outcome and propensity models, $\hat{\mu}$ and $\hat{\delta}$. In general, these models can be obtained through parametric or nonparametric regression methods. Appealing properties of the estimator arise from its second-order nuisance estimation error, that is,
@@ 2.2 An augmented doubly robust estimator $\hat{\bm{\beta}}_{UW}$
Under the same conditions in Theorem 2.2, further assume that both of the nuisance parameters are consistent, and $\|(1-{\delta_{i}}/{\hat{\delta_{i}}})(\hat{\nu_{i}}-\nu_{i})\|_{2}=o_{\mathbb{P}}(n^{-1/2})$. Then the estimator $\hat{\bm{\beta}}_{UW}$ defined in (4) is asymptotically normal:
@@
$\hat{\Sigma}_{UW}=\left(\frac{1}{n}\sum_{i=1}^{n}\mathbf{W}_{i}\mathbf{W}_{i}^{T}\right)^{-1}\left(\frac{1}{n}\sum_{i=1}^{n}(\hat{Y}_{i}^{UW}-\mathbf{W}_{i}^{T}\hat{\bm{\beta}}_{UW})^{2}\mathbf{W}_{i}\mathbf{W}_{i}^{T}\right)\left(\frac{1}{n}\sum_{i=1}^{n}\mathbf{W}_{i}\mathbf{W}_{i}^{T}\right)^{-1}$ (5)
@@
The goal of this analysis is to identify peptides whose abundances are marginally associated with a given feature, such as disease status, cell types, or other phenotypic variables, which is a commonly studied problem in the field. Therefore, we adopt a regression model $Y^{j}=\mathbf{W}^{T}\beta^{j}+\epsilon^{j}$ for peptide $j$, which doesn’t include any auxiliary peptide data. For imputation, we use other peptide data as auxiliary variables because we want to approximate $Y^{j}$ with less variability, and we assume that the noise variables $\epsilon^{1},\cdots,\epsilon^{p}$ are correlated between peptides. In this way, we examine the marginal association between the variables of interest while incorporating other peptide data outside the model without contradiction.
Remark 2.6.
@@ 3.1 The input data and the algorithm
For $n$ i.i.d. samples, the observed abundances of $p$ peptides can be written as an $n\times p$ matrix $\mathbf{C}\odot\mathbf{Y}\in\mathbb{R}^{n\times p}$, where $\mathbf{C}\in\{0,1\}^{n\times p}$ indicates the entry-wise missingness and $\mathbf{Y}\in\mathbb{R}^{n\times p}$ is the full data matrix without missing. Here $\odot$” stands for the component-wise product. Only $\mathbf{C}$ and $\mathbf{C}\odot\mathbf{Y}$ are available. Also observed is a covariate data matrix $\mathbf{W}\in\mathbb{R}^{n\times q}$. Note that, for each observation $i$, the observed abundance of peptides $\{Y_{ij}:C_{ij}=1,j\in[p]\}$ provide extra information to impute the unobserved ones $\{Y_{ij}:C_{ij}=0,j\in[p]\}$. Consequently, we treat $\mathbf{U}_{i}^{j}:=\{C_{i\ell}Y_{i\ell}:\ell\neq j\}$ as the augmented covariate for each peptide $j$. The inference task is to test the significance of regression coefficients for the low dimensional covariates in $\mathbf{W}$ on each peptide. To this end, we will obtain individual p-values for each peptide using the asymptotic results presented in the previous section and then apply a multiple-testing framework such as the Benjamini-Hochberg procedure (Benjamini and Hochberg, 1995).
The procedure involves the estimation of two nuisance functions: the propensity score function $\delta$ and the augmented regression function $\nu$. In our upcoming experiments, we use Logistic regression to estimate $\delta$. The estimation of $\nu$ requires repeatedly regressing each column of $\mathbf{Y}$ on both the low-dimensional covariate $\mathbf{W}$ and the other peptides as the high-dimensional auxiliary covariate $\mathbf{U}$. An additional challenge is that each column of $\mathbf{Y}$ has many missing entries, even when used as a covariate in the regression problem, resulting in unregular auxiliary covariates for different samples. To address this issue, we use a nonparametric deep neural network model, VAEIT (Du et al., 2022, see Appendixl B for details), which allows for flexible input and simultaneous estimation of the multi-response regression. More specifically, VAEIT utilizes masking strategies to model the mapping $\nu_{j}:\mathbf{W}_{i},\mathbf{U}_{i}^{j}\mapsto Y_{ij}$ for $j\in[p]$. In other words, it learns regression functions $\nu_{j}(\mathbf{W}_{i},\mathbf{U}_{i}^{j})=\mathbb{E}[Y_{ij}\mid\mathbf{W}_{i},\mathbf{U}_{i}^{j}]$ for $j\in[p]$, enabling the simultaneous imputation of all missing peptides. Although the convergence rates of estimating individual nuisance functions may be slower than $\sqrt{n}$, the doubly robust procedure allows for valid statistical inference as long as the product of the two convergence rates is $o_{\mathbb{P}}(n^{-1/2})$, as illustrated in Theorem 2.3. Finally, the whole procedure is summarized in Algorithm 1.
@@
To achieve the double robustness advantages, the proposed method requires consistent estimates of the nuisance parameters. The consistency of the propensity score estimator $\hat{\delta}$ follows through theoretical arguments. The logistic model is nonlinear in parameters and some iterative procedures, such as Newton-type methods, are applied to find the MLE of parameters. When the number of parameters is far less than the sample size, then such procedures are known to achieve a global convergence (Gourieroux and Monfort, 1981; Lee, Sun and Saunders, 2014). Rashid and Shifa (2009) provide an extensive Monte Carlo simulation result to verify a finite sample performance. The consistency of $\hat{\nu}$ is more difficult to prove, and we provide an empirical result to show that the mean squared error consistently decreases as the sample size increases in the simulated dataset (Figure C1).